# Answer, Assemble, Ace: Understanding How Transformers Answer Multiple Choice Questions

### v1

@@ 4.3 Initial Observations
Vocabulary projection reveals patterns distinct to MCQA.
Since vocabulary projection can produce small values in earlier layers of a network simply because hidden states have not yet been transformed into a space mappable to high probabilities on individual tokens, we first compare results on output hidden states of individual layers between the formatted MCQA task on Colors, and a greedy generation version, where we provide instances in the format ‘‘Corn is yellow. What color is corn?’’ and take the first greedy-decoded token as the prediction (i.e., we expect the model to produce‘‘ yellow’’).666All models investigated achieve 100% accuracy on this task given 3 in-context examples. Results are shown in Fig. 3.
We observe, notably, that while the greedy version of the task and the “top token” line demonstrate that hidden states can be transformed by the unembedding matrix into non-negligible probabilities as early as the output of layer 22 (for Olmo; layer 18 for Llama2), non-negligible probabilities on answer symbols emerge a couple of layers later (layer 23 for both models). Additionally, from these noted layers onward, the relative difference between the predicted token and the second choice is much smaller than in the greedy version. We conclude that it takes some extra layers of processing before output hidden states decisively predict an answer choice symbol (~layer 30 for Olmo and layer 32 for Llama2), even though updates to hidden state representations after layer 22/18 (Olmo/Llama2) are generally not changing models’ greedy-generated color predictions.
Correctness does not dictate predictive patterns.
We next investigate a dataset where a model performing consistent MCQA (Olmo-7B-SFT) does not achieve perfect performance: HellaSwag. We separate predictions by two attributes: correctness of the prediction, and which letter is predicted (‘‘A’’ or ‘‘B’’). Fig. 11 (Appendix) reveals that vocabulary projection does not reveal any meta-notions of correctness, as trends in vocabulary space align more with what token will be predicted (diagonal pairs) as opposed to whether that token is correct (vertical pairs). The same trend holds for activation patching (Fig. 12, Appendix). Thus, as we are most interested in correct predictions, we will only include correctly-predicted instances in subsequent plots (i.e., Figs. 5(a) and 5(b) for vocabulary projection and Fig. 6(a) for activation patching).
@@
As elucidated in Fig. 6(a), activation patching highlights the key causal role that layer 21 plays in encoding the answer choice that Olmo-7B-SFT will predict. While layers prior to 21 may promote or demote a specific token choice, their contributions to the residual stream are overridden by layer 21’s, and similarly, later layers appear to be carrying forward information contributed to the residual stream by layer 21. In Llama2-13B-Chat (Fig. 16(a)), we observe a more diffuse window of ~6 layers (15-21) at which activation patching is increasingly able to change the model’s prediction.
Compared with Fig. 6(a), Figs. 5(a) and 5(b) illustrate the complementary nature of the two methods: activation patching reveals important mechanisms before hidden states are projectable to the vocabulary space, while vocabulary projection shows how relevant information in the residual stream appears in the vocabulary space over remaining layers. Indeed, one hypothesis that arises from these results is that it takes a couple of layers of processing for information encoded in a hidden state to become mapped to higher probabilities on those tokens.
@@ 5 Roles of Specific Layers
Key Layers are consistent across datasets.
We observe projected token probabilities for Olmo-7B-SFT across all three datasets in Fig. 4, and activation patches for two datasets in Figs. 6(a) and 6(b). Despite Olmo-7B-SFT’s varying performance on the datasets, key layers are highly consistent. Notably, layer 21 also plays a key causal role in the promotion of answer symbol tokens for Colors. Similar results are found for Llama2-13B-Chat (Figs. 14, 16(a) and 16(b)). These findings corroborate the hypothesis that the same mechanisms are responsible for promoting answer symbol letters, regardless of task complexity.
@@
When using a different symbol space (i.e., ‘‘C:_D:_’’), processing trends are noticeably different for the Olmo model (Figs. 5(c) and 5(d)). The model first assigns probability to the labels that are more likely or expected (‘‘A’’ and ‘‘B’’), before making an abrupt switch to the correct symbol (‘‘C’’ or ‘‘D’’) at layer 30. Activation patching reveals the same trend (Figs. 6(c) and 6(d)): patching in prompts that promote the production of ‘‘C’’ and ‘‘D’’ causes the relative difference between ‘‘A’’ and ‘‘B’’ to close at layers 21-29, before an abrupt change to predicting ‘‘C’’ driven by layer 30. This provides evidence that Olmo-7B-SFT solidifies final label predictions for “OOD” prompts in later layers, and could explain why some models struggle with OOD formats.777Note that this is specific to Olmo-7B-SFT. Llama2-13B-Chat immediately promotes ‘‘C’’ and ‘‘D’’ (Figs. 16(c), 16(d), 17(c) and 17(d)).
@@ 6 Roles of Layer Components
We next investigate the role that MHSA and MLP functions play in the observed behavior.
@@
In Figs. 7(a) and 7(b), we directly project the hidden states output by these functions (positions A and B in Fig. 10). Large spikes in probabilities come from both MHSA (layers 23, 30, 31) and MLP outputs (layers 28, 30). The breakdown in spiking behavior by attention heads in Fig. 8(a) (bottom) shows just how sparse the number of attention heads projecting to non-negligible probability values is (~8 total), even in layers where the MHSA output hidden state assigns large probabilities to ‘‘A’’ and ‘‘B’’ (recall that the MHSA output vector is a sum of these weighted attention head vectors). Activation patching results (Figs. 7(c) and 7(d)) further elucidate that the MHSA mechanism is driving the encoding of relevant information to predict the answer at layer 21, and at layer 29 for ‘‘C:_D:_’’ (Fig. 18, Appendix). The importance of the MHSA function as the source of the causal effect observed at layers 15-21 is also evident for Llama2-13B-Chat (Fig. 20, Appendix), though more attention heads in the model project to the vocabulary tokens (Fig. 21(a)).888It is unclear why vocabulary projections of the two functions are so different from activation patches for the Llama2 model.
Individual attention heads dominate the production of answer choice symbols (over MLPs).
Plotting all attention heads and MLP layers that assign high probability to the sum of answer choices () in Fig. 8(a) (top; Fig. 21(a) for Llama2-13B-Chat), we see that attention heads generally dominate the distribution, despite the fact that they only contain ~half the learnable parameters of MLPs (~67M vs. ~135M per layer). However, attention heads have access to global context which an MLP does not have, making them especially useful for resolving back-references to portions of the input sequence or copying information from one position to another.
The same network components that promote the production of any answer choice promote specific answer choices as well.
Increasing probability on both answer choices is not the same as being discriminative, i.e., promoting one choice at the expense of another. We plot all heads and MLPs that are discriminative () in Figs. 8(b) and 8(c) (top; Figs. 21(b) and 21(c) for Llama2-13B-Chat). There is 100% overlap: every plotted component also appears in Fig. 8(a).
Model components specialize.
We observe in Figs. 8(b) and 8(c) (bottom) attention heads that take on different roles depending on the predicted symbol. Of the 13 total model components in the two plots, we observe:
1.
3 that consistently assign high probability to ‘‘A’’. While they are consistently assigning higher probability to ‘‘A’’ than to ‘‘B’’, their average magnitudes drop when ‘‘B’’ is predicted.
2.
4 that consistently assign high probability to ‘‘B’’. While they consistently assign higher probability to ‘‘B’’ than to ‘‘A’’, their average magnitudes drop when ‘‘A’’ is predicted.
3.
3 that assign high probability to whichever token that is ultimately predicted, analogous to “correct letter heads” Lieberum et al. (2023).
4.
3 that only activate when ‘‘B’’ is predicted, though all assigning probabilities .
@@ 7 Where are poorly-performing models going wrong?
In Fig. 9 (Table 3 for other models), our synthetic task confirms that there is a crucial point at which formatted MCQA skill is learned for Olmo-v1.7-7B: between 350B and 1T tokens. Performance goes from near-random to near-100%, even for checkpoints with poor performance on more challenging datasets.999As a reminder, all models achieve 100% accuracy on the color string generation version of the task (Section 4.3), meaning that low performance can be attributed to formatted MCQA. This result highlights the value of a synthetic task: it helps to disentangle to what extent poor MCQA performance on a dataset is due to a lack of dataset-specific knowledge vs. an inability to perform formatted MCQA.
@@
We use vocabulary projection to inspect differences between these checkpoints on the Colors task. Fig. 22 (Appendix) illustrates the separability between the top and second-top answer choices, which grows as a function of training for more steps. Indeed, the primary distinction observed between the strong checkpoints (Figs. 22(b) and 22(c)) and the weak checkpoint (Fig. 22(a)) is the separability of ‘‘A’’ and ‘‘B’’, particularly on and after layer 29. This is in spite of the fact that the 350B checkpoint model predicts ‘‘B’’ as the final answer on 99.5% of instances (Table 3), and may explain why small edits to network parameters can decrease label bias Li and Gao (2024). Finally, we observe in Fig. 23 (Appendix) that Olmo v1.7 7B’s ability to do symbol binding as demonstrated on Colors does not alleviate the token separability problem on other datasets on which the model is less performant. Overall, these results demonstrate that an inability to disentangle answer symbol tokens in vocabulary space, not an inability to assign high probability to these tokens, is a property of poorly-performing models.
@@ 8 Conclusion
This work provides novel insights into the mechanisms language models employ when answering formatted multiple-choice questions. Our analysis leverages two complementary techniques: vocabulary projection and activation patching. Using carefully-designed prompts and a synthetic task, we uncover a series of stages, attibutable to a very sparse set of specialized network components, in which transformers select, promote, and align (in the case of OOD prompts) the answer choice symbol they predict. With MCQA datasets such as MMLU routinely used as a performance signal for model development, it is important to disentangle and test formatted MCQA ability in order to build models that operate in more robust and reliable ways. Our work makes an important first step towards this goal.
@@ Acknowledgments
We thank members of the Aristo team at AI2 and the H2Lab at the University of Washington for feedback and discussions.
Limitations
Activation patching makes some assumptions about the independence of model components to make computation tractable (namely, that one can patch in individual hidden states in isolation to measure their effect on the network as opposed to patching all possible combinations of states). A nascent line of work focuses on making activation patching more efficient so that a larger number of states and/or state combinations can be intervened on Kramár et al. (2024); Syed et al. (2023); this is an interesting direction for future work. We additionally only patch on pairs of prompts where the model predicted both labels correctly; it could be interesting to compare behavior on consistently incorrect instances instead.
Finally, we primarily focus on a couple of performant small-to-medium-size models in this work, but expanding to other model families (and sizes) capable of formatted MCQA is an interesting direction for future work in order to understand the generality of our findings.
@@ Appendix A Appendix
A.1 Plot Complexity w.r.t Number of Answer Choices
The extent to which a model component prefers ‘‘A’’ over other answer choices in a binary setup is equivalent to ; this quantity simultaneously illustrates (in the negative direction) the extent to which the component prefers ‘‘B’’. However, for 3 answer choices, say ‘‘A’’, ‘‘B’’, ‘‘C’’, 3 quantities are needed to measure preference for each choice: , , and .
A.2 Dataset Details
HellaSwag (Zellers et al., 2019) is a 4-way multiple-choice commonsense natural language inference dataset. The goal is to select the best completion for a given sentence prompt. The test set is not publicly available; we sample a fixed set of 5000 instances from the validation set used in our experiments (apart from activation patching, for which we sample 1000 instances due to compute restrictions). We use the first 3 training set instances as in-context examples for all instances.
MMLU Hendrycks et al. (2021), or the “Massive Multitask Language Understanding” benchmark, spans 57 different topical areas. The questions are 4-way multiple-choice spanning subjects in social sciences, STEM, and humanities that were manually scraped from practice materials available online for exams such as the GRE and the U.S. Medical Licensing Exam. We sample a fixed set of 5000 instances from the test set used in our experiments (apart from activation patching, for which we sample 1000 instances due to compute restrictions). We sample a fixed set of 3 in-context example instances from the 5 provided for each topical area.
We convert both tasks to binary multiple-choice following the procedure in Section 3.
A.3 Transformer Background
The output hidden states of transformer models Vaswani et al. (2017) are a linear combination of the outputs of non-linear functions (i.e., multi-head self-attention and multi-layer perceptrons) at each layer. For the Olmo and Llama2 models, which sequentially apply multi-head self-attention (MHSA) and a multi-layer perceptron (MLP), for input hidden state , the output hidden state of layer is defined by the following equation:
Where and represent the multi-head self-attention operation and multi-layer perceptron, respectively, defined by their unique parameters at layer , and LN represents layer normalization.101010 also receives as input the hidden states at layer from other token positions in the rollout (as dictated by the attention mask), which we omit here for brevity. Note that represents simple vector addition, which serves to establish residual connections. This is possible because the output of each operation (MHSA, MLP, and LN) is a vector . We visualize a single layer’s block structure in Fig. 10.
@@ A.3 Transformer Background
The output of a MHSA function can be further broken down into a sum of the output of each attention head. For the input vector to layer , , the output of MHSA is computed as the concatenation of each head’s vector output, , times an output weight matrix which can be simplified into a sum as follows (originally elucidated in Elhage et al. (2021)):
@@
| Model | |||||||||
| Olmo-7b | Olmo-v1.7-7b | Llama2-7b | Llama2-13b | ||||||
| Prompt | Base | SFT | Base@350B | Base@1T | Base | Base | Chat | Base | Chat |
| ‘‘A:_B:_’’ | |||||||||
| ‘‘A:_B:_’’ | |||||||||
[…]
| ‘‘C:_D:_’’ | |||||||||
| ‘‘C:_D:_’’ | |||||||||
[…]
| Model | |||||||||
| Olmo-7b | Olmo-v1.7-7b | Llama2-7b | Llama2-13b | ||||||
| Prompt | Base | SFT | Base@350B | Base@1T | Base | Base | Chat | Base | Chat |
| ‘‘A:_B:_’’ | |||||||||
| ‘‘A:_B:_’’ | |||||||||
[…]
| ‘‘C:_D:_’’ | |||||||||
| ‘‘C:_D:_’’ | |||||||||
[…]
| Model | |||||||||
| Olmo-7b | Olmo-v1.7-7b | Llama2-7b | Llama2-13b | ||||||
| Prompt | Base | SFT | Base@350B | Base@1T | Base | Base | Chat | Base | Chat |
| ‘‘A:_B:_’’ | |||||||||
| ‘‘A:_B:_’’ | |||||||||
[…]
| ‘‘C:_D:_’’ | |||||||||
| ‘‘C:_D:_’’ | |||||||||
[…]
| Model | ||||||||||
| Olmo-7b | Olmo-v1.7-7b | Llama2-7b | Llama2-13b | |||||||
| Prompt | Statistic | Base | SFT | Base@350B | Base@1T | Base | Base | Chat | Base | Chat |
[…]
| vs. ‘‘A:_B:_’’ | Consistently Incorrect | |||||||||
[…]
| Incons. Correct (Both ‘‘B’’) | ||||||||||
[…]
| vs. ‘‘C:_D:_’’ | Consistently Incorrect | |||||||||
[…]
| Model | ||||||||||
| Olmo-7b | Olmo-v1.7-7b | Llama2-7b | Llama2-13b | |||||||
| Prompt | Statistic | Base | SFT | Base@350B | Base@1T | Base | Base | Chat | Base | Chat |
[…]
| vs. ‘‘A:_B:_’’ | Consistently Incorrect | |||||||||
[…]
| vs. ‘‘C:_D:_’’ | Consistently Incorrect | |||||||||
[…]

### v2

@@ 4.4 Initial Observations
A/B/C/DA/B/C/D
A/B/C/DA/B/C/D (Colors)
Q/Z/R/X A/B/C/D (HellaSwag)
1/2/3/4 A/B/C/D (HellaSwag)
@@
As elucidated in Fig. 3(b), activation patching highlights the key causal role that layer 24 plays in encoding the answer choice that Olmo 7B Instruct will predict on Colors. While layers prior to 24 may promote or demote a specific token choice, the predicted token does not change, indicating that these layers’ contributions to the residual stream are either nonexistent, or are overridden by layer 24’s. Similarly, later layers appear to be carrying forward information contributed to the residual stream by layer 24 (while also slightly reducing the logits of the other answer choices), since they do not have any effect at undoing the change in predicted token caused by intervening on layer 24’s output representation.
Compared with Fig. 3(b), Fig. 4(a) illustrates the complementary nature of the two methods: activation patching reveals important mechanisms before hidden states are projectable to the vocabulary space, while vocabulary projection shows how relevant information in the residual stream appears in the vocabulary space over remaining layers. Indeed, one hypothesis that arises from these results is that it takes a couple of layers of processing for information encoded in a hidden state to become mapped to higher scores on those tokens.
@@ 5 Roles of Specific Layers
Key Layers differ across models and # of shots, but are consistent across datasets and predicted tokens.
Fig. 4 presents projected token scores for Olmo 7B Instruct across all three datasets, and Figs. 3(a) and 3(b) activation patches for two datasets. Despite Olmo 7B Instruct’s varying performance on the datasets, key layers are largely consistent. Notably, layer 24 also plays the key causal role in the promotion of answer symbol tokens for HellaSwag, though layers 22-23 also have a small effect. In Fig. 4, the outputs of layers 26 and 29 lead to the largest logit and probit increases for the predicted answer token, with layer 31 serving to boost the logits of all tokens (this has no effect in probability space). Similar results are found for the Llama and Qwen models (Figs. 14 and 15), corroborating the hypothesis that the same mechanisms are responsible for promoting answer symbol letters across tasks, regardless of task complexity. However, comparing these graphs with Fig. 4 reveals that logit magnitudes and key layers differ across models, even for those of comparable size.
Comparing Fig. 4 (3-shot) to 0-shot results (Fig. 16), the role of in-context examples primarily manifests in not only increasing the logit values of all answer choices in the last layers of the model (29-31), but also leading to a substantial increase in the predicted token’s value such that it far surpasses other tokens in the vocabulary and accumulates nearly all the probability mass. This does not happen in the 0-shot case despite overall accuracy being similar, and also coincides with a noticeable (temporary) demotion of valid answer symbols at layer 28.
We break down Fig. 4(b) by predicted answer letter (Fig. 5), observing that trends are generally similar across A, B, C, and D for Olmo 7B Instruct and Llama 3.1 8B Instruct (Fig. 18).888We observe similar results in activation patching experiments to Figs. 3(a) and 3(b) when using different correct answer letters, further corroborating this trend. Qwen 2.5 1.5B exhibits a fairly strong early preference for A at layer 20 which may be due to the token’s corpus frequency, but is otherwise similar across symbols (Fig. 19), particularly from layer 22 onward.
@@
Processing trends are noticeably different for some models when using a different symbol space (Q/Z/R/X, 1/2/3/4, or another set of random letter symbols, O/E/B/P). For random letters, Olmo and Qwen models first assign non-negligible probability to labels that are more likely or expected (A, B, C, D) even though they are not included in the prompt, before making an abrupt switch to the correct symbols (Q/Z/R/X or O/E/B/P) at layers 29 and 23, respectively (Fig. 6, top, Fig. 22 and Fig. 24). Activation patching reveals the same trend (Figs. 3(c) and 17). For example, for the Olmo model, patching in from prompts containing Q/Z/R/X or O/E/B/P only decisively promotes the new symbol from layer 29, noticeably later than layer 24 in the A/B/C/D experiments (Figs. 3(a) and 3(b)). This provides evidence that Olmo 7B Instruct solidifies final label predictions for “OOD” prompts in later layers, and could explain why some models struggle with OOD formats. In the case of a more standard prompt, 1/2/3/4, scores assigned to A/B/C/D remain negligible (Fig. 6, bottom).999It is of note that activation patching from A/B/C/D to 1/2/3/4 (Fig. 3(d)) also identifies layer 29 as a key layer. This differs from Fig. 3(a), where position but not symbols changed. Since correct answer position does not change in this experiment, one possible explanation is that layer 24 is causally attributed to encoding positional information, while layer 29 is attributed to encoding the relevant answer token at that position.
@@ 6 Roles of Layer Components
We next investigate the role that MHSA and MLP functions play in the observed behavior.101010Vocabulary projections in logit space of a layer’s MLP and MHSA functions are a direct additive decomposition of the projections on the layer’s final hidden state. This is also true for individual attention heads weighted by their respective rows of the MHSA output matrix, since the MHSA output vector is a weighted sum of the heads’ output vectors. See Eqs. 1, 2 and 3 in Appendix A.2 for the derivation.
Self-attention mechanisms dominate the production of answer choice symbols (over MLPs).
In Figs. 7(a) and 7(b), we directly project the hidden states output by these functions (positions A and B in Fig. 12). Both functions map to large increases in logit values at layer 29. However, MLP outputs serve to increase the logit value of all answer choices, not to be discriminative (as noted by the near-0 logit difference line). Increases in the logit difference are attributed to the MHSA functions at layers 26, 27, and 29. This is despite the fact that MHSA functions only contain ~half the learnable parameters of MLPs (~67M vs. ~135M per layer). However, attention heads have access to global context which an MLP does not have, making them especially useful for resolving back-references to portions of the input sequence or copying information from one position to another.
@@
The corresponding breakdown in the logit and probit differences from projecting individual attention head output vectors to the vocabulary space (Fig. 8(a)) shows that this effect is quite sparse, with 1-4 attention heads per layer (out of 32) projecting to non-negligible values on any answer choice symbol. We additionally observe high absolute magnitude overlap between Fig. 8(a) and Fig. 8(b)/8(c), indicating that the same components promoting probability on any answer choices (Fig. 8(a)) are also generally discriminative in preferring one choice over the other (Fig. 8(b)/8(c)). However, the differences between Figs. 8(b) and 8(c) reveal that some heads play letter-specific roles. Similar trends hold for the Qwen model (Fig. 26). Activation patching results (Figs. 7(c) and 7(d)) further elucidate that the MHSA mechanism is driving the encoding of relevant information to predict the answer primarily at layer 24, with Fig. 13 demonstrating that a single attention head is responsible for this effect.
when A correct
when A correct
when B correct
@@ 7 Where are poorly-performing models going wrong?
In Fig. 9, our synthetic task confirms that there is a crucial point at which formatted MCQA skill is learned for Olmo 0724 7B base – between 80k and 100k training steps. Performance goes from near-random to near-100%, even for checkpoints with poor performance on more challenging datasets. This result highlights the value of a synthetic task: it helps to disentangle to what extent poor MCQA performance on a dataset is due to a lack of dataset-specific knowledge vs. an inability to perform formatted MCQA.
@@
We use vocabulary projection to inspect differences between these checkpoints on the Colors task. Fig. 27 (Appendix) illustrates the separability between the answer choice symbols, which grows as a function of training for more steps. While at and steps the model represents a meaningful ranking of the answer tokens that results in high accuracy using the multiple-choice prediction rule, only with sustained training does the model widen the logit difference substantially, leading to the model assigning high probability to the predicted answer alone. The small differences in logit scores observed at earlier checkpoints may explain why small edits to network parameters are effective at decreasing label bias (Li & Gao, 2024).
Finally, we observe in Fig. 28 (Appendix) that the ability to do symbol binding at the final training checkpoint, as demonstrated on Colors, does not alleviate small logit differences between answer symbols, even for correct predictions, on datasets on which the model is less performant: scores assigned to correct answers are both higher variance and tempered in layers 30-31 in these cases.
@@ 8 Conclusion
This work provides novel insights into the mechanisms language models employ when answering formatted multiple-choice questions. Our analysis leverages two complementary techniques: vocabulary projection and activation patching. Using carefully designed prompts and a synthetic task, we uncover a series of stages, attributable to a sparse set of specialized network components, in which models select, promote, and align (for OOD prompts) the answer choice symbol they predict. With MCQA datasets such as MMLU routinely used as a performance signal for model development, it is important to disentangle and test formatted MCQA ability in order to build models that operate in more robust and reliable ways. Our work makes an important first step towards this goal.
Reproducibility Statement
To the best of our ability, we have included all details necessary to reproduce our experiments in the text of this paper and accompanying Appendix. We have also open-sourced our code (https://github.com/allenai/understanding_mcqa) and Colors dataset (https://huggingface.co/datasets/sarahwie/copycolors_mcqa). We fixed random seeds to ensure full reproducibility of our experiments.
@@ Acknowledgments
We thank the anonymous reviewers, members of the Aristo team at AI2, and members of the H2Lab at the University of Washington for valuable feedback. YB was supported by the Israel Science Foundation (grant No. 448/20), an Azrieli Foundation Early Career Faculty Fellowship, and an AI Alignment grant from Open Philanthropy. This research was partially funded by the European Union (ERC, Control-LM, 101165402). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Appendix
A.1 Dataset Details
HellaSwag (Zellers et al., 2019) is a 4-way multiple-choice commonsense natural language inference dataset. The goal is to select the best completion for a given sentence prompt. We sample a fixed set of 1000 instances from the test set used in our experiments. We sample 3 random training set instances to serve as in-context examples.
MMLU (Hendrycks et al., 2021), or the “Massive Multitask Language Understanding” benchmark, spans 57 different topical areas. The questions are 4-way multiple-choice spanning subjects in social sciences, STEM, and humanities that were manually scraped from practice materials available online for exams such as the GRE and the U.S. Medical Licensing Exam. We sample a fixed set of 1000 instances from the test set used in our experiments. We sample a fixed set of 3 in-context example instances from the 5 provided for each topical area.
In-context example selection.
We randomly select the correct answer position for each in-context example (which comes to positions 012; or A B C, Q R X, or 1 2 3, depending on the prompt format. We find that models have more consistent performance when in-context examples are used (particularly, the models we benchmark which are not instruction-tuned), but we do also include 0-shot results for a subset of strong models, where the inclusion of in-context examples does not have a large effect on performance (Fig. 12).
A.2 Transformer Background
The output hidden states of transformer models (Vaswani et al., 2017) are a linear combination of the outputs of non-linear functions (i.e., multi-head self-attention and multi-layer perceptrons) at each layer. For most models, which sequentially apply multi-head self-attention (MHSA) and a multi-layer perceptron (MLP), for input hidden state , the output hidden state of layer is defined by the following equation:
Where and represent the multi-head self-attention operation and multi-layer perceptron, respectively, defined by their unique parameters at layer , and LN represents layer normalization.111111 also receives as input the hidden states at layer from other token positions in the rollout (as dictated by the attention mask), which we omit here for brevity. Note that represents simple vector addition, which serves to establish residual connections. This is possible because the output of each operation (MHSA, MLP, and LN) is a vector . We visualize a single layer’s block structure in Fig. 12.
@@ A.2 Transformer Background
The output of a MHSA function can be further broken down into a sum of the output of each attention head. For the input vector to layer , , the output of MHSA is computed as the concatenation of each head’s vector output, , times an output weight matrix which can be simplified into a sum as follows (originally elucidated in (Elhage et al., 2021)):
@@
Appendix B Strengths and Weaknesses of Vocabulary Projection vs. Activation Patching
Activation patching makes some assumptions about the independence of model components to make computation tractable (namely, that one can patch in individual hidden states in isolation to measure their effect on the network as opposed to patching all possible combinations of states). A nascent line of work focuses on making activation patching more efficient so that a larger number of states and/or state combinations can be intervened on Kramár et al. (2024); Syed et al. (2024); this is an interesting direction for future work.
(Llama 3.1 8B Instruct)
(Qwen 2.5 1.5B Instruct)
(Olmo 7B 0724 Instruct)
(% acc.)
(% acc.)
(% acc.)
(% acc.)